OpenStudy (anonymous):

Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. 6x + 2 + 3x = 10x + 2 – x

OpenStudy (radar):

9x+2=9x+2 This is an identity with infinite solutions.

OpenStudy (anonymous):

6x + 2 + 3x = 10x + 2 - x 9x + 2 = 9x + 2 {combined like terms} 0 = 0 {subtracted 9x and subtracted 2 from both sides} = infinite solutions ?